# Compressible and immiscible fluids with arbitrary density ratios: diffuse interface theory

### v3

@@ VII.2 Equation of the state (EOS)
which is the EOS of ideal gas with a factor, $\ln\rho$. Note that the thermodynamic pressure $\Psi$ should be used instead of $p$. The term $\ln\rho$ is analogous to the entropy contribution, and this result is consistent with previous work on the free energy formulation of a water-gas system Ledesma-Aguilar et al. (2014). By dividing both sides of Eq. (97) by the molar volume, we obtain $\Psi\sim\rho\ln\rho$, which is the prescribed EOS used in Refs. Ledesma-Aguilar et al. (2014); Ridl and Wagner (2018). This highlights our central result: the density evolution in our theory naturally generates the EOS, in contrast to the prior EOS imposed in existing diffuse-interface methods.
@@ VIII Conclusion
$\mathbf{u}\dot{\rho}=-\nabla\cdot[(p\mathbf{I}+\underline{\underline{\bm{\Theta}}})]-\rho\nabla\cdot[\tau_{1}(\nabla\bm{\zeta}+\nabla\bm{\zeta}^{T})+\tau_{2}(\nabla\cdot\bm{\zeta})\mathbf{I}]$ (112)
@@
  • the local pressure $p$ and stress tensor $\underline{\underline{\bm{\Theta}}}$;
@@ Acknowledgements.
This research is supported by VirtMat project of the Helmholtz association, as part of the programme “MSE-materials science and engineering” number-43.31.01. F.W. conceived the overall concept, conducted all theoretical derivations, and prepared the manuscript. B.N. contributed to securing the funding.
@@ Appendix C Derivative of $\mu\phi$
Supposing that $\phi(x)$ and $\phi(y)$ are functions: $\mathbb{R}^{3}\mapsto\mathbb{R}$, and $\mu$ is a functional: $\Phi\mapsto\mathbb{R}$, where $\phi(x),\ \phi(y)\in\Phi$, the derivative for the product of $\mu$ and $\phi$ reads Gelfand et al. (2000)
$\frac{\delta[\mu\phi(x)]}{\delta\phi(y)}=\phi(x)\frac{\delta\mu}{\delta\phi(y)}+\mu\frac{\delta\phi(x)}{\delta\phi(y)}.$ (120)
@@
In a special case, $\phi(x)=\phi(y)$, we have $\frac{\delta\phi(x)}{\delta\phi(y)}=1$, and therefore $\delta(\mu\phi)=\phi\delta\mu+\mu\delta\phi$. As $\mu=\delta e/\delta\phi$, the above equation is finally expressed as $\delta(\mu\phi)=\phi\delta\mu+\delta e$.

### v4

@@ VII.2 Equation of the state (EOS)
which is the EOS of ideal gas with a factor, $\ln\rho$. Note that the thermodynamic pressure $\Psi$ should be used instead of $p$. The term $\ln\rho$ is analogous to the entropy contribution, and this result is consistent with previous work on the free energy formulation of a water-gas system Ledesma-Aguilar et al. (2014). By dividing both sides of Eq. (97) by the molar volume, we obtain $\Psi\sim\rho\ln\rho$, which is the prescribed EOS used in Refs. Ledesma-Aguilar et al. (2014); Ridl and Wagner (2018). This highlights our central result: the density evolution in our theory naturally generates EOS, in contrast to a prescribed EOS in existing diffuse-interface methods.
@@ VIII Conclusion
$\mathbf{u}\dot{\rho}=-\nabla\cdot[(p\mathbf{I}+\underline{\underline{\boldsymbol{\Theta}}})]-\rho\nabla\cdot[\tau_{1}(\nabla\boldsymbol{\zeta}+\nabla\boldsymbol{\zeta}^{T})+\tau_{2}(\nabla\cdot\boldsymbol{\zeta})\mathbf{I}]$ (112)
@@
  • the local pressure $p$ and stress tensor $\underline{\underline{\boldsymbol{\Theta}}}$;
@@ Acknowledgements.
We express our sincere gratitude to the Editor, Dr. Cottin-Bizonne, for facilitating constructive exchanges with the referees, and to all referees for their careful and insightful comments.
@@ Appendix C Derivative of $\mu\phi$
Supposing that $\phi(\mathbf{x})$ and $\phi(\mathbf{y})$ are functions: $\mathbb{R}^{3}\mapsto\mathbb{R}$, and $\mu$ is a functional: $\Phi\mapsto\mathbb{R}$, where $\phi(\mathbf{x}),\ \phi(\mathbf{y})\in\Phi$, the derivative for the product of $\mu$ and $\phi$ reads Gelfand et al. (2000)
$\frac{\delta[\mu\phi(\mathbf{x})]}{\delta\phi(\mathbf{y})}=\phi(\mathbf{x})\frac{\delta\mu}{\delta\phi(\mathbf{y})}+\mu\frac{\delta\phi(\mathbf{x})}{\delta\phi(\mathbf{y})}.$ (120)
@@
In a special case, $\phi(\mathbf{x})=\phi(\mathbf{y})$, we have $\frac{\delta\phi(\mathbf{x})}{\delta\phi(\mathbf{y})}=1$, and therefore $\delta(\mu\phi)=\phi\delta\mu+\mu\delta\phi$. As $\mu=\delta e/\delta\phi$, the above equation is finally expressed as $\delta(\mu\phi)=\phi\delta\mu+\delta e$.
Appendix D The problem of the classic continuity equation for compressible fluids
For one-component fluids, the mass conservation reads
$\frac{\partial}{\partial t}\int_{V}nm_{0}\mathrm{d}V=-\int_{\Gamma}\mathbf{j}\cdot\mathbf{n}\mathrm{d}\Gamma,$ (121)
where $n$ is the number of the mass particles, such as molecules or atoms, and $m_{0}$ is the mass per particle. The flux $\mathbf{j}$ represents the flow of mass particles across the boundary $\Gamma$, with normal vector $\mathbf{n}$. By applying Gauss’s divergence theorem, we obtain
$\frac{\partial(nm_{0})}{\partial t}+\nabla\cdot\mathbf{j}=0.$ (122)
To convert Eq. (122) into an evolution equation for the density, both sides must be divided by the volume element $V$. Based on the following calculation for $V$:
$V=v_{m}+v_{e},$ (123)
where $v_{m}$ is the volume of molecules or atoms, and $v_{e}$ is the excess volume which can also be interpreted as the vacancy volume, we classify the discussion into two cases: (I) When the vacancy volume is much smaller than the molecular volume, i.e. i.e., $v_{e}\ll v_{m}$, as in a one-component solid or a solid-like liquid, the system corresponds to a nearly incompressible phase. (II) When the vacancy volume is comparable to the molecular volume, as in a one-component gas, the system corresponds to a compressible fluid. In case (I), the system cannot be significantly compressed, and the local volume element $V$ can be considered nearly constant. By defining the flux $\mathbf{j}=m_{0}n\mathbf{u}$ and multiplying both sides by the approximately constant factor $1/V\approx 1/v_{m}$, we recover the classic continuity equation
$\frac{\partial\rho}{\partial t}+\nabla\cdot(\rho\mathbf{u})=0.$ (124)
In case (II), the local volume element $V$ can vary with time and space. Consequently, the continuity equation, Eq. (124) takes the form:
$(nm_{0})\frac{\partial(1/V)}{\partial t}+\frac{1}{V}\frac{\partial(nm_{0})}{\partial t}+\nabla\cdot(\frac{nm_{0}}{V}\mathbf{u})=0.$ (125)
Due to the large amount of vacancies in case II, $(1/V)$ is not guaranteed to be a conserved variable, since vacancies can appear or disappear locally under the influence of pressure. As such, conventional models require a prescribed equation of state (EOS) to define $1/V$, for example, $pV=\text{constant}$ for an ideal gas. Moreover, when $(1/V)$ is not conserved, the convection flux cannot be formulated as $\mathbf{j}=\rho\mathbf{u}$, because the density $\rho$ contains information from $1/V$. Incorporating a non-conserved variable into the mass conservation equation is therefore not appropriate.
The same issue arises for multicomponent compressible fluids with $K\in\mathbb{N}$ components, where $(1/V)$ is also not conserved. The only difference is that the volume element is rewritten as $V=\sum_{i=1}^{K}v_{i}+v_{e}$, where $v_{i}$ is the volume of the $i$-th component.
In the static state with $\mathbf{u}=\mathbf{0}$, Eq. (124) or Eq. (125) returns to a trivial equality: $0=0$, which has no physical meaning. In this case, we have to consider the microscopic velocity to derive EOS (Section VII B). Within a moving interface indicating that $1/V$ is non-uniform in space, the continuity equation Eq. (124) cannot be applied, as $\partial_{t}(1/V)+\nabla\cdot(\mathbf{u}/V)$ is not necessary to be zero, since the vacancy can vanish or appear locally in the moving process of the interface.
In our model, we consider two distinct evolution equations. The first is for the volume concentration $\phi$ including the excess volume and the volumes of all other mass species. The evolution of $1/V$ is driven by the maximization of the mixing entropy $s_{\phi}$, resulting in a diffusion equation. For single-component fluids, this diffusion occurs via the exchange of molecules with vacancies, such as in self-diffusion. For multicomponent fluids, diffusion occurs through the exchange of molecules with both vacancies and other molecular species. The second evolution equation governs the density. For the approximation of inviscid fluids, the density evolution in our model based on Eq. (75) is expressed as
$\mathbf{u}\frac{\mathrm{d}\rho}{\mathrm{d}t}=\mathbf{u}\bigg(\frac{\partial\rho}{\partial t}+\mathbf{u}\cdot\nabla\rho\bigg)=-\nabla p.$ (126)
A heuristic understanding for Eq. (126) is that $\nabla p\propto\nabla(1/V)$ under a prescribed EOS, which accounts for local compression effects on the volume element and thereby extends the classical continuity equation. For incompressible fluids, the pressure work is nearly zero, $pdV\approx 0$, and therefore $\nabla p\approx 0$ in the Gibbs-Duhem relation. In this case, Eq. (126) reduces to the classic continuity equation for incompressible flow. This scenario corresponds to $v_{e}=0$ i.e., a rigorously incompressible fluid with no vacancies. For multicomponent incompressible fluids, the density evolution of each component $\rho_{i}$ can be solved separately. For compressible fluids ($v_{e}\neq 0$), the density evolution of individual components cannot generally be solved separately, since the effect of $v_{e}$ on $v_{1}$, $v_{2}$, $\cdots$, $v_{K}$ is unknown. In our model for compressible fluids, Eq. (126) naturally gives rise to the equation of state (Section VII B), so a prescribed EOS is not required. However, if a prescribed EOS is available Ridl and Wagner (2018), which implicitly prescribes the effect of $v_{e}$, then the density evolution of different components can, of course, be solved separately.